OpenStudy (anonymous):

A storm in Sydney, Australia, caused $30 million in damage. That estimate was in Australian dollars. At that time, 1.5 Australian dollars were worth 1 American dollar. Determine the number of American dollars of damage the storm caused.

OpenStudy (a_clan):

AUD 1.5 = USD 1 AUD 1 = USD 1/1.5 AUD 30m= USD (1/1.5)*30m = USD 20m